1 point) Suppose X is a normally distributed random variable with H9 and ơ 1.3. Find each of the following probabilities: (a) P(12 < X < 15) (b) P(6.1 K X K 16.7)- (c) P(11.1 K X K 16.7)- (d) P(X 2 11.5)- (e) P(X s 16.7)-
Answer #1

Solution :

Given that,

mean = \mu = 9

standard deviation = \sigma =1.3

a ) P (12 \leq x \leq 15)

P (12 - 9 / 1.3)  \leq ( x -  \mu/ \sigma)\leq ( 15- 9 / 1.3)

P ( 3 / 1.3 \leq z \leq 6 / 1.3 )

P ( 2.31 \leq z \leq 4.62 )

P (z \leq 4.62) - p ( z \leq 2.31 )

Using z table

= 1- 0.9896

= 0.0104

Probability = 0.0104

b ) P (6.1\leq  x \leq 16.7 )

P (6.1 - 9 / 1.3 ) \leq  ( x -  \mu/ \sigma)\leq  (16.7 - 9 / 1.3)

P ( -2.9/1.3\leq z \leq 7.7/1.3)

P ( - 2.23\leq  z \leq 5.92 )

P (z \leq  5.92)p ( z \leq-2.23 )

Using z table

= 1- 0.0129

= 0.9871

Probability =0.9871

c) P (11.1\leq  x \leq 16.7 )

P (11.1-9/ 1.3 ) \leq  ( x -  \mu/ \sigma)\leq  (16.7 - 9 / 1.3)

P (2.1/3\leq z \leq 7.7/1.3)

P (1.62\leq  z \leq 5.92 )

P (z \leq  5.92)p ( z \leq1.62 )

Using z table

= 1- 0.9474

= 0.0526

Probability =0.0526

d) P ( x \geq  11.5 )

= 1 -P( x \leq 11.5 )

= 1 -P ( x -  \mu/ \sigma)\leq  (11.5 - 9 / 1.3)

= 1 -P( z \leq 2.5/1.3)

= 1 -P(z \leq 1.92 )

Using z table

= 1- 0.9726

= 0.0274

Probability =0.0274

e) P ( x \leq 16.7 )

P ( x -  \mu/ \sigma)\leq  (16.7 - 9 / 1.3)

P (z \leq 7.7/1.3)

P ( z \leq 5.92 )

Using z table

= 1

Probability =1